OpenStudy (anonymous):

\[n\in \mathbb{N}\] Show that : \[\frac{1}{\sqrt{n}}-\frac{1}{\sqrt{n+1}}=\frac{1}{n\sqrt{n+1}+(n+1)\sqrt{n}}\] Tipp: expand with \[(\frac{\sqrt{n+1}+\sqrt{n}}{\sqrt{n+1}+\sqrt{n}})\]

OpenStudy (zarkon):

hard to prove something that isn't true

OpenStudy (anonymous):

i am sorry, i wrote wrong

OpenStudy (zarkon):

take the left hand side and make it into one fraction. Next, multiply it by your 'tip' and simplify

OpenStudy (anonymous):

i found left side \[\frac{1}{2n+1+2n\sqrt{1}}\] but i dont know somewhat it feels wrong

OpenStudy (zarkon):

\[\frac{1}{\sqrt{n}}-\frac{1}{\sqrt{n+1}}\] \[=\frac{\sqrt{n+1}-\sqrt{n}}{\sqrt{n}\sqrt{n+1}}\]

OpenStudy (zarkon):

\[=\frac{\sqrt{n+1}-\sqrt{n}}{\sqrt{n}\sqrt{n+1}}\times\frac{\sqrt{n+1}+\sqrt{n}}{\sqrt{n+1}+\sqrt{n}}\] \[=\frac{(n+1)-(n)}{\sqrt{n}\sqrt{n+1}\sqrt{n+1}+\sqrt{n}\sqrt{n+1}\sqrt{n}}\] \[=\frac{1}{\sqrt{n}(n+1)+n\sqrt{n+1}}\]

OpenStudy (zarkon):

which is equivalent to what you have above (using commutativity)

OpenStudy (anonymous):

thx Zarkon, now i get it